Let a and b be integers. Prove that (a, b) = 1 if and only if (a + b, ab) = 1.
1
Expert's answer
2019-10-08T01:51:04-0400

1)Suppose that "(a,b)=d>1" . So "a" and "b" are divisible by "d". Then "a+b" and "ab" are divisible by "d", that is "(a+b,ab)" is divisible by "d" .

So if "(a+b,ab)=1", then "(a,b)=1".

2)Suppose that "(a+b,ab)" is divisible by some prime "p". Then "ab" and "a+b" are divisible by "p".

Since "a+b" is divisible by "p", there is such "k" that "a+b=pk".

Since "ab" is divisible by "p", we have that "a" is divisible by "p" or "b" is divisible by "p".

If "a" is divisible by "p", then "a=pl" for some "l", then "b=pk-a=p(k-l)", that is "b" is divisible by "p".

By symmetry we obtain that if "b" is divisible by "p", then "a" is divisible by "p".

In both cases "a" and "b" are divisible by "p", so "(a,b)" is divisible by "p".

Therefore if "(a,b)=1", then "(a+b,ab)=1".

From 1) and 2) we obtain that "(a+b,ab)=1" if and only if "(a,b)=1".